Give the Trigonometric Functions and the value of x and y components



1. 30 degrees, West of South


2. 60 degrees North if West

1
Expert's answer
2021-11-22T10:12:56-0500

Suppose we have a vector "\vec A" .


1. "A_x=|\vec A|\cdot\sin30°" and "A_y=|\vec A|\cdot\cos30°"


2. "A_x=|\vec A|\cdot\cos60°" and "A_y=|\vec A|\cdot\sin60°"